A new record for running the stairs of the Empire State Building was set on February 3, 2003. The 86 flights, with a total of 1576 steps, was run in 9 minutes and 33 seconds.

Part A If the height gain of each step was 0.20 m, and the mass of the runner was 71 kg , what was his average power output during the climb? Give your answer in both watts and horsepower.

Part B Express your answer using two significant figures. P = m hp
Answer #1

Part A :

No.of steps n= 1576

Height of each step = 0.20 m

Total height H = 1576 x 0.20

= 315.2 m

Workdone w = mgh

= 71 x 9.81 x 315.2

= 219539.952 J

Output Power P = w/t

time taken t = 9min33s

= 573 s

Output Power P =  219539.952 / 573

= 383.141277 watts

= 383.141277/746 hp

= 0.51359 hp

PartB :

P =0.51359 hp

= 0.51 hp